OpenStudy (anonymous):

What volume of 0.10 M NaOH is required to neutralize 30 mL of 0.10 M HCl? Does anybody know a formula that would help me find this?

OpenStudy (anonymous):

@Destinymasha @dan815

OpenStudy (anonymous):

First you have to set up the equation: \[NaOH + HCl \rightarrow H_2O + NaCl\]

OpenStudy (anonymous):

And then you just have to do a bunch of conversions: \[30 mL(\frac{ 1L }{ 1000mL })(\frac{ 0.1molHCl }{ 1L })(\frac{ 1 mol NaOH }{ 1 mol HCl })(\frac{ 1L }{ 0.1molNaOH })(\frac{ 1000mL }{ 1L })\]

OpenStudy (anonymous):

Thanks! so the answer would be 30 mL?